# Many-body critical phase in a quasiperiodic chain and dynamical Widom lines in Fock space properties

### v1

@@ IV.1 Noninteracting systems
In addition to the self-energy, we also calculate the averaged FS IPR as shown in Fig. 6(a-d) for the same transitions shown in Fig. 5(a-d). Since, here we deal with noninteracting many-body eigenstate which can be written using single particle eigenfunctions as |$\ket{\Psi}=\prod_{\nu=1}^{N}c_{\nu}^{\dagger}\ket{0}$ where $c_{\nu}=\sum_{i}\psi_{\nu}(i)c_{i}$ are the diagonalizing operator of the single particle Hamiltonian $H_{sp}$, the IPR can be written as $IPR=\sum_{I}|\braket{I}{{\Psi}}|^{4}/(\sum_{I}|\braket{I}{{\Psi}}|^{2})^{2}$, where $\braket{I}{{\Psi}}$’s are given by Slater determinants. This way we can avoid the exact diagonalization of the many body Hamiltonian and study the noninteracting system in FS for larger system sizes up to $L=28$. We choose the energy window and number of disorder realizations such that we average over at least a total of $5000$ eigenstates. The Widom-like line $\lambda=2$ is manifested as weak dip in IPR in the critical phase [Fig. 6(c)] and the Widom line $\mu=1$ appears as a broad peak in IPR in the localized phase [Fig. 6(d)].
@@
[…]
@@
We extract the coefficient $A_{I}$ and the fractal dimension $D_{2}$ from the $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$ through a linear fitting of the data shown in Fig. 7(a,c,e). From this analysis we conclude that $D_{2}\approx 1$, $A_{I}<1$ in the delocalized phase, whereas $10<D_{2}<$ 1, $A_{I}>1$ in the localized phase. The critical phase exhibits IPR in between the delocalized and localized phases, namely $10<D_{2}<$ 1 like the localized phase, but $A_{I}>1$ as in the delocalized phase. We also extract similar coefficients and exponents, $A_{s}$ and $D_{s}$, respectively, from $\Delta_{t}$, as summarized in Table 1 (see also Appendix B). However, $\Delta_{t}$ is rather irregular as a function of $\mathcal{N}_{F}$ in the critical phase, unlike the IPR [Fig. 7(c)] and, as a result, we could not extract $A_{s}$ and $D_{s}$ reliably in the critical phase.
@@ IV.2 Interacting systems
[…]
Figure 8: Fock-space self-energy for interacting systems: Typical values of self-energy $\Delta_{t}$ across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$ In figures (c) and (d), the vertical dashed lines denote the non-monotonic dependence of $\Delta_{t}$ and IPR on the parameters in MBC and MBL phases, respectively. The solid vertical lines denote the points of the phase transition.
[…]
Figure 9: Fock-space $IPR$ for interacting systems: $IPR$ across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$. In figures (c) and (d), the vertical dashed lines denote the peak/dip of $\Delta_{t}$ across the Widom lines in Fig.1(b). The solid vertical lines denote the phase transitions.
[…]
@@
Scaling ansatz Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025) similar to Eq.(6) can be written for the FS self-energy $\Delta_{t}$. Using these scaling ansatz for $\Delta_{t}$, we have also performed the finite-size scaling analysis for the data of self-energy across ergodic-MBL, ergodic-MBC and MBC-MBL transitions in Figs.8(a,b,c,d). The analysis across the ergodic-MBL transition, as discussed in Appendix B, leads to qualitatively similar results as in the earlier studies on random and other quasiperiodic systemsSutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025). However, due to more irregularity and non-monotonicity of $\Delta_{t}(\mathcal{N}_{F},\lambda,\mu)$ compared to that in $IPR$ in the MBC phase, we could not obtain reliable data collapse across the transitions to MBC phase. Hence, we have only discussed the finite-size scaling analysis of $IPR$ above.
IV.3 Distributions of the imaginary part of self-energy and inverse participation ratio
[…]
Figure 13: Distributions of $IPR$: (a-c) Probability distributions of $IPR$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$ and MBC $(\mu=1.5,\lambda=1.0)$ phases, respectively. (d) The same in the delocalized $(\mu=0.5,\lambda=1.0)$ phase of the non-interacting EAAH model.
@@ V Fock-space localization length
In this section, we analyze the off-diagonal elements $G_{IJ}(E=0)$ of the FS propagator Roy et al. (2023); Ghosh et al. (2025) in the ergodic, MBL and MBC phases of the interacting EAAH model, and compare them with the phases of the non-interacting model. We calculate the geometric mean or the typical value, defined as $G(r_{IJ})=\text{exp}[\langle\ln G_{IJ}\rangle]$, where delimited-⟨$\langle$..$\rangle$ … ⟩ denotes an average over $\phi$ and all the off-diagonal elements $G_{IJ}$ for a pair of FS sites are at a hopping distance $r_{IJ}$. The latter is the minimum number of nearest-neighbor hops to go from $I$ to $J$ in the middle slice. The number of realizations used for $\phi$ are $2000,1000,400,200,100$ for $L=12,14,16,18,20$, respectively. Plots of $\ln[G(r_{IJ})]$ as a function of $r_{IJ}$ are shown in Figs. 14(a-c) and Figs. 14(d-f) for the phases of the interacting and non-interacting EAAH model. In all the phases, the plots show a linear regime with negative slope, i.e., $\ln[G(r_{IJ})]\propto-r_{IJ}$, before it deviates from linearity depending on $L$. This deviation from the linear regime corresponds to rare hopping distances and associated multiple length scales in FS Roy et al. (2023); Ghosh et al. (2025). The linear regime implies the existence of a decay length $\xi_{F}(L)$, dubbed as the ‘FS localization length’ Roy et al. (2023); Ghosh et al. (2025), such that $G(r_{IJ})\sim\text{exp}[-r_{IJ}/\xi_{F}(L)]$.
[…]
@@
[…]
@@
VI Discussion and Conclusions
@@
In the thermodynamic limit, the NEE phases, like the MBC phase, is expected to be more fragile than the MBL phase. The stability of the MBL phase in the thermodynamic limit for systems with random disorder has been called to question due to the possibility of long-range resonances and avalanche instabilities from rare weak-disorder regions in larger system sizes, not accessible via exact diagonalization $(L\leq 24)$ Sels and Polkovnikov (2021); Sels (2022); Crowley and Chandran (2022); Long et al. (2023); Morningstar et al. (2022); Sierant and Zakrzewski (2022). However, unlike the systems with random disorder, the quasiperiodic systems are not susceptible to the usual avalanche instability De Roeck and Huveneers (2017) due to the absence of rare weak-disorder regions Agarwal et al. (2017). Hence the MBC phases, like the MBL phase, in the quasiperiodic EAAH model could be more robust. The mechanism of obtaining NEE states from the single-particle critical states is a new one and may produce more stable NEE states than found in previous works Li et al. (2015); Deng et al. (2017) on the quasiperiodic GAAH model.
@@ Acknowledgements
SB acknowledges support from CRG, SERB (ANRF), DST, India (File No. CRG/2022/001062) and STARS, MoE, Govt. of India (File. No. MoE-STARS/STARS-2/2023-0716) .
@@ Appendix A Energy level-spacings and spectral form factor
We discuss here the analysis of distributions $P(s)$ of consecutive many-body energy level-spacings $s$, normalized by the (arithmetic) mean level spacing, in the ergodic, MBL and MBC phases of the interacting EAAH model. The nearest neighbor level-spacing distributions for ergodic, MBL and MBC phases are shown in Fig. 16(a-c), respectively, for increasing $L$. For the largest system size $L=16$, we fit the data to the Brody distribution given by $P(s)=As^{a}\exp(-As^{a+1}/(a+1))$ where $A=(a+1)\Gamma(\frac{a+2}{a+1})^{(a+1)}$. $a=1$ and $a=0$ correspond to GOE and Poisson distributions, respectively. We find $a\approx 0.94$, $0.02$ and $0.26$ for the choices of parameters in Fig. 16(a-c), respectively. The plots indicate that $P(s)$ approaches GOE distribution in the ergodic phase and Poisson distribution in the MBL phase Atas et al. (2013). Moreover, for the MBC phase $0.260.15<a<$ 0.26, implying distributions intermediate between GOE and Poisson, where maximum value of $a\approx 0.26$ appears along the Widom line ($\lambda\approx 2$) in the MBC phase.
[…]
Figure 16: Energy level statistics and correlations in the interacting EAAH model: (a-c) Level spacing distributions $P(s)$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$, MBC $(\mu=1.5,\lambda=2.0)$ phases, respectively. The dashed lines are fit to the Brody distribution to the data for $L=16$ (d-f) Spectral form factor ($SFF$) of the interacting systems for the same parameters as in figures (a-c), respectyively. For calculation of $SFF$, the number of mid-spectrum states $N_{st}=1000$.
@@
where delimited-⟨$\langle$..$\rangle$ … ⟩ stands for an average over some ensemble, e.g., realizations of $\phi$ in the EAAH model. At $\tau=0$, $SFF=N_{st}^{2}$. For very small $\tau\ll\tau_{D}$ compared to the Thouless time $\tau_{D}$, $\langle Z^{*}(\tau)Z(\tau)\rangle=\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$ due to absence of spectral correlation and $SFF$ decreases showing non-universal model-specific spectral features. At $\tau=\tau_{H}$, the Heisenberg time, $\tau$ becomes comparable to inverse of mean level spacing. As a result, the second term in Eq. 7 vanishes on average and a plateau at $SFF=N_{st}$ appears. For $\tau_{D}<\tau<\tau_{H}$, $SFF$ shows a ramp indicating the development of spectral correlation such that $\langle Z^{*}(\tau)Z(\tau)\rangle\neq\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$. This region shows universal features and does not appear in the absence of any spectral correlation. In Fig. 16(d-f) rescaled $SFF/N_{st}$ is plotted for the three phases corresponding to Fig. 16(a-c), respectively. The ergodic phase shows a sharper and longer ramp [Fig. 16(a)], whereas MBC phase shows a weaker and shorter ramp [Fig. 16(c)], indicating the presence of long-range correlations in the spectrum for both ergodic and MBC phases. In the MBL phase, no ramp regime is found indicating absence of spectral correlation [Fig. 16(b)] Prakash et al. (2021).
@@ Appendix B Fock-space self energy and finite-size scaling
[…]
@@
Here we first make an attempt to extract $A_{s}$ and $D_{s}$ from $\Delta_{t}=A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ for delocalized-localized and ergodic-MBL transitions in non-interacting and interacting systems, respectively. $A_{s}$ and $D_{s}$, extracted through a linear fitting to the plots in Fig.Fig. 17(a) [also Fig. 5(a)], are shown in Fig. 17(b) across the delocalized-localized transition as a function of $\lambda$ for $\mu=0.5$ in the the non-interacting EAAH model. We find ($D_{s}\approx 1$, $A_{s}<1$) and ($10<D_{s}<$ 1, $A_{s}>1$) deep in the delocalized and localized phases, respectively.

### v2

@@ IV.1 Noninteracting systems
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)           Refer to caption(e)           Refer to caption(f)
@@
We extract the coefficient $A_{I}$ and the fractal dimension $D_{2}$ from the $IPR\simeq A_{I}\mathcal{N}_{F}^{-D_{2}}$ through a linear fitting of the data shown in Fig. 7(a,c,e). From this analysis we conclude that $D_{2}\approx 1$, $A_{I}<1$ in the delocalized phase, whereas $0<D_{2}<1$, $A_{I}>1$ in the localized phase. The critical phase exhibits IPR in between the delocalized and localized phases, namely $0<D_{2}<1$ like the localized phase, but $A_{I}>1$ as in the delocalized phase. We also extract similar coefficients and exponents, $A_{s}$ and $D_{s}$, respectively, from $\Delta_{t}$, as summarized in Table 1 (see also Appendix B). However, $\Delta_{t}$ is rather irregular as a function of $\mathcal{N}_{F}$ in the critical phase, unlike the IPR [Fig. 7(c)] and, as a result, we could not extract $A_{s}$ and $D_{s}$ reliably in the critical phase.
@@ IV.2 Interacting systems
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 8: Fock-space self-energy for interacting systems: Typical values of self-energy $\Delta_{t}$ across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$ . In figures (c) and (d), the vertical dashed lines denote the non-monotonic dependence of $\Delta_{t}$ and IPR on the parameters in MBC and MBL phases, respectively. The solid vertical lines denote the points of the phase transition.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 9: Fock-space IPRIPR for interacting systems: IPRIPR across (a) the ergodic-MBL transition as a function of $\lambda$ for $\mu=0.5$, (b) ergodic-MBC transition as a function of $\mu$ for $\lambda=1.0$, (c) MBC-MBL transitions as a function of $\lambda$ for $\mu=1.5$, and (d) MBL-MBC transition as a function of $\mu$ for $\lambda=3.5$, for increasing system sizes $L$. In figures (c) and (d), the vertical dashed lines denote the peak/dip of $\Delta_{t}$ across the Widom lines in Fig.1(b). The solid vertical lines denote the phase transitions.
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)           Refer to caption(e)           Refer to caption(f)
@@
Scaling ansatz Sutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025) similar to Eq.(6) can be written for the FS self-energy $\Delta_{t}$. Using these scaling ansatz for $\Delta_{t}$, we have also performed the finite-size scaling analysis for the data of self-energy across ergodic-MBL, ergodic-MBC and MBC-MBL transitions in Figs.8(a,b,c,d). The analysis across the ergodic-MBL transition, as discussed in Appendix B, leads to qualitatively similar results as in the earlier studies on random and other quasiperiodic systemsSutradhar et al. (2022); Roy et al. (2023); Ghosh et al. (2025). However, due to more irregularity and non-monotonicity of $\Delta_{t}(\mathcal{N}_{F},\lambda,\mu)$ compared to that in $IPR$ in the MBC phase, we could not obtain reliable data collapse across the transitions to MBC phase. Hence, we have only discussed the finite-size scaling analysis of $IPR$ above. The distributions of IPR and $\Delta_{I}$ for different phases show distinct features which are discussed in Appendix C.
@@ V Fock-space localization length
In this section, we analyze the off-diagonal elements $G_{IJ}(E=0)$ of the FS propagator Roy et al. (2023); Ghosh et al. (2025) in the ergodic, MBL and MBC phases of the interacting EAAH model, and compare them with the phases of the non-interacting model. We calculate the geometric mean or the typical value, defined as $G(r_{IJ})=\text{exp}[\langle\ln G_{IJ}\rangle]$, where $\langle$$\rangle$ denotes an average over $\phi$ and all the off-diagonal elements $G_{IJ}$ for a pair of FS sites are at a hopping distance $r_{IJ}$. The latter is the minimum number of nearest-neighbor hops to go from $I$ to $J$ in the middle slice. The number of realizations used for $\phi$ are $2000,1000,400,200,100$ for $L=12,14,16,18,20$, respectively. Plots of $\ln[G(r_{IJ})]$ as a function of $r_{IJ}$ are shown in Figs. 12(a-c) and Figs. 12(d-f) for the phases of the interacting and non-interacting EAAH model. In all the phases, the plots show a linear regime with negative slope, i.e., $\ln[G(r_{IJ})]\propto-r_{IJ}$, before it deviates from linearity depending on $L$. This deviation from the linear regime corresponds to rare hopping distances and associated multiple length scales in FS Roy et al. (2023); Ghosh et al. (2025). The linear regime implies the existence of a decay length $\xi_{F}(L)$, dubbed as the ‘FS localization length’ Roy et al. (2023); Ghosh et al. (2025), such that $G(r_{IJ})\sim\text{exp}[-r_{IJ}/\xi_{F}(L)]$.
     Refer to caption(a)           Refer to caption(d)           Refer to caption(b)           Refer to caption(e)           Refer to caption(c)           Refer to caption(f)
@@
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
@@
VI Widom-like line in the presence of random potential
In this short section, we make an attempt to find Widom-like line in presence of random disorder. It is not clear at the moment how to find a Widom-like line in a truly random system as appearance of Widom line would typically need presence of a special critical line or a triple point. However, one can think of a single particle model with both a regular hopping $(J=1)$ and quasiperidoc hopping (of strength $\mu$), like in EAAH model plus a random onsite potential (of strength $W$) instead of the quasiperiodic potential. In that case, for $W=0$ there will be a transition from delocalized to critical at $\mu=1$, according to Fig. 1(a) in the main text. For finite non-zero $W$ all single particle states become localized for all values of $\mu$. Interestingly, we find that the states show the strongest localization along the $\mu=1$ line for $W>0$. Hence, one can say that a Widom-like line emanates from $(\mu=1,W=0)$ point and continues for finite $W$ (see Fig. 14). Presumably this $\mu=1$ line will continue to persist even in the stipulated MBL phase in presence of interaction.
     Refer to caption(a)           Refer to caption(b)
Figure 14: Tight-binding model with quasiperiodic hopping and random onsite potential: (a) For $W\neq 0$, $IPR$ as a function of $\mu$ shows a peak at around $\mu=1$ tracing the presence of Widom-like line. (b) Phase diagram: In absence of onsite random potential $(W=0)$, the single particle system undergoes delocalized-to-critical phase transition at $\mu=1$ denoted by the black dot. In presence of random potential $(W\neq 0)$, all the single particle eigenstates become localized. However, a Widom-like line denoted by the red dot-dashed line emanates from the critical point $\mu=1$ showing a line of the strongest localization as demonstrated in figure (a).
VII Discussion and Conclusions
@@
In the thermodynamic limit, the NEE phases, like the MBC phase, is expected to be more fragile than the MBL phase. The stability of the MBL phase in the thermodynamic limit for systems with random disorder has been called to question due to the possibility of long-range resonances and avalanche instabilities from rare weak-disorder regions in larger system sizes, not accessible via exact diagonalization $(L\leq 24)$ Sels and Polkovnikov (2021); Sels (2022); Crowley and Chandran (2022); Long et al. (2023); Morningstar et al. (2022); Sierant and Zakrzewski (2022). However, unlike the systems with random disorder, the quasiperiodic systems are not susceptible to the usual avalanche instability De Roeck and Huveneers (2017) due to the absence of rare weak-disorder regions Agarwal et al. (2017). Hence the MBC phases, like the MBL phase, in the quasiperiodic EAAH model could be more robust. The mechanism of obtaining NEE states from the single-particle critical states is a new one and may produce more stable NEE states than found in previous works Li et al. (2015); Deng et al. (2017) on the quasiperiodic GAAH model. Nevertheless, the issue with extrapolating the existence of MBC phase, based on the small finite-size numerics, to the thermodynamic limit is expected to be more complex than that for the MBL phase. In this regard, it might be useful to apply matrix-product state (MPS) or density-matrix renormalization group based approaches for the EAAH model to access larger system sizes, as has been attempted Pomata et al. (2023) for the GAAH model. However, the MBC states have volume-law entanglement and thus harder to describe through MPS-based approaches than the MBL phase.
@@ Acknowledgements
SB acknowledges support from CRG, SERB (ANRF), DST, India (File No. CRG/2022/001062) and STARS, MoE, Govt. of India (File. No. MoE-STARS/STARS-2/2023-0716)
@@ Appendix A Energy level-spacings and spectral form factor
We discuss here the analysis of distributions $P(s)$ of consecutive many-body energy level-spacings $s$, normalized by the (arithmetic) mean level spacing, in the ergodic, MBL and MBC phases of the interacting EAAH model. The nearest neighbor level-spacing distributions for ergodic, MBL and MBC phases are shown in Fig. 15(a-c), respectively, for increasing $L$. For the largest system size $L=16$, we fit the data to the Brody distribution given by $P(s)=As^{a}\exp(-As^{a+1}/(a+1))$ where $A=(a+1)\Gamma(\frac{a+2}{a+1})^{(a+1)}$. $a=1$ and $a=0$ correspond to GOE and Poisson distributions, respectively. We find $a\approx 0.94$, $0.02$ and $0.26$ for the choices of parameters in Fig. 15(a-c), respectively. The plots indicate that $P(s)$ approaches GOE distribution in the ergodic phase and Poisson distribution in the MBL phase Atas et al. (2013). Moreover, for the MBC phase $0.15<a<0.26$, implying distributions intermediate between GOE and Poisson, where maximum value of $a\approx 0.26$ appears along the Widom line ($\lambda\approx 2$) in the MBC phase.
     Refer to caption(a)           Refer to caption(d)           Refer to caption(b)           Refer to caption(e)           Refer to caption(c)           Refer to caption(f)
Figure 15: Energy level statistics and correlations in the interacting EAAH model: (a-c) Level spacing distributions $P(s)$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$, MBC $(\mu=1.5,\lambda=2.0)$ phases, respectively. The dashed lines are fit to the Brody distribution to the data for $L=16$ . (d-f) Spectral form factor (SFFSFF) of the interacting systems for the same parameters as in figures (a-c), respectyively. For calculation of $SFF$, the number of mid-spectrum states $N_{st}=1000$.
@@
where $\langle$$\rangle$ stands for an average over some ensemble, e.g., realizations of $\phi$ in the EAAH model. At $\tau=0$, $SFF=N_{st}^{2}$. For very small $\tau\ll\tau_{D}$ compared to the Thouless time $\tau_{D}$, $\langle Z^{*}(\tau)Z(\tau)\rangle=\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$ due to absence of spectral correlation and $SFF$ decreases showing non-universal model-specific spectral features. At $\tau=\tau_{H}$, the Heisenberg time, $\tau$ becomes comparable to inverse of mean level spacing. As a result, the second term in Eq. 7 vanishes on average and a plateau at $SFF=N_{st}$ appears. For $\tau_{D}<\tau<\tau_{H}$, $SFF$ shows a ramp indicating the development of spectral correlation such that $\langle Z^{*}(\tau)Z(\tau)\rangle\neq\langle Z^{*}(\tau)\rangle\langle Z(\tau)\rangle$. This region shows universal features and does not appear in the absence of any spectral correlation. In Fig. 15(d-f) rescaled $SFF/N_{st}$ is plotted for the three phases corresponding to Fig. 15(a-c), respectively. The ergodic phase shows a sharper and longer ramp [Fig. 15(a)], whereas MBC phase shows a weaker and shorter ramp [Fig. 15(c)], indicating the presence of long-range correlations in the spectrum for both ergodic and MBC phases. In the MBL phase, no ramp regime is found indicating absence of spectral correlation [Fig. 15(b)] Prakash et al. (2021).
@@ Appendix B Fock-space self energy and finite-size scaling
        Refer to caption(a)                Refer to caption(b)                 Refer to caption(c)                 Refer to caption(d)
@@
Here we first make an attempt to extract $A_{s}$ and $D_{s}$ from $\Delta_{t}=A_{s}\mathcal{N}_{F}^{-1+D_{s}}$ for delocalized-localized and ergodic-MBL transitions in non-interacting and interacting systems, respectively. $A_{s}$ and $D_{s}$, extracted through a linear fitting to the plots in Fig.Fig. 16(a) [also Fig. 5(a)], are shown in Fig. 16(b) across the delocalized-localized transition as a function of $\lambda$ for $\mu=0.5$ in the the non-interacting EAAH model. We find ($D_{s}\approx 1$, $A_{s}<1$) and ($0<D_{s}<1$, $A_{s}>1$) deep in the delocalized and localized phases, respectively.
@@
Appendix C Distributions of the imaginary part of self-energy and inverse participation ratio
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
     Refer to caption(a)           Refer to caption(b)           Refer to caption(c)           Refer to caption(d)
Figure 19: Distributions of IPRIPR: (a-c) Probability distributions of $IPR$ in the ergodic $(\mu=0.5,\lambda=1.0)$, MBL $(\mu=0.5,\lambda=5.0)$ and MBC $(\mu=1.5,\lambda=1.0)$ phases, respectively. (d) The same in the delocalized $(\mu=0.5,\lambda=1.0)$ phase of the non-interacting EAAH model.